OpenStudy (anonymous):

2. Solve for y, then find the value of y when given x = –2. 6x = 7 – 4y (1 point) –12 24 3. Solve for h. A = bh (1 point) = h 2A = h = Ab h = 2Ab

OpenStudy (anonymous):

6x=7-4y where x=-2 6(-2)=7-4y plug in negative 2 -12=7-4y subtract 7 from both sides -7 =-7 -19=-4y divide by negative 4 on both sides 19/4=y or if you need mixed numbers that would be y=4& 3/4